OpenStudy (anonymous):

Solve by using the perfect squares method. x2 + 8x + 16 = 0 A. 2 B. –2 C. –4 D. 4

OpenStudy (zzr0ck3r):

\(x^2+bx+c=0\implies (x+\frac{b}{2})^2=-c+(\frac{b}{2})^2\) solve for x

OpenStudy (zzr0ck3r):

so \((x+4)^2=-16+4^2\) solve for x

OpenStudy (zzr0ck3r):

or note \((a+b)^2=a^2+2ab+b^2\) for you \(a=x, b=4\)